A number is 12 more than the other. Find the numbers if their sum is 48.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the first number be x and the second number be x + 12

As per the given question;

Sum of the two numbers = 48

x + x + 12 = 48

2x + 12 = 48

2x = 48 – 12

2x = 36

x = 36/2

x = 18


So the answer is 18 and 30.


Thanks